Part 1

Next day Aladdin invited the Sultan to see the palace. On entering the hall with the four-and-twenty windows with their rubies, diamonds and emeralds, he cried, “It is a world’s wonder! There is only one thing that surprises me. Was it by accident that one window was left unfinished?” “No, sir, by design,” returned Aladdin. “I wished your Majesty to have the glory of finishing this palace.” The Sultan was pleased, and sent for the best jewelers in the city. He showed them the unfinished window, and bade them fit it up like the others. “Sir,” replied their spokesman, “we cannot find jewels enough.” The Sultan had his own fetched, which they soon used, but to no purpose, for in a month’s time the work was not half done. Aladdin knowing that their task was vain, bade them undo their work and carry the jewels back, and the genie finished the window at his command. The Sultan was surprised to receive his jewels again, and visited Aladdin, who showed him the window finished. The Sultan embraced him, the envious vizier meanwhile hinting that it was the work of enchantment.

Part 2

Aladdin had won the hearts of the people by his gentle bearing. He was made captain of the Sultan’s armies, and won several battles for him, but remained as courteous as before, and lived thus in peace and content for several years.

But far away in Africa the magician remembered Aladdin, and by his magic arts discovered that Aladdin, instead of perishing miserably in the cave, had escaped, and had married a princess, with whom he was living in great honour and wealth. He knew that the poor tailor’s son could only have accomplished this by means of the lamp, and travelled night and day till he reached the capital of China, bent on Aladdin’s ruin. As he passed through the town he heard people talking everywhere about a marvelous palace. “Forgive my ignorance,” he asked, “what is the palace you speak of?” Have you not heard of Prince Aladdin’s palace,” was the reply, “the greatest wonder in the world? I will direct you if you have a mind to see it.” The magician thanked him who spoke, and having seen the palace knew that it had been raised by the Genie of the Lamp, and became half mad with rage. He determined to get hold of the lamp, and again plunge Aladdin into the deepest poverty.